# Accurate Pocket Identification for Binding-Site-Agnostic Docking

### v1

@@ VII Evaluation on the Astex Diverse Set
Figure 22: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 37, and Kalasanty 38 for the Astex Diverse Set 29.
Figure 23: Docking accuracy of AutoDock Vina 12 for the Astex Diverse Set 29 when guided by different pocket prediction algorithms, comparing Top-1 and ensemble accuracy.
Figure 24: Distribution of search grids leading to poses that pass the RMSD $<$ 2 Å test, for Top-1 Vina poses and for subleading poses when the Top-1 Vina pose fails, for Astex Diverse Set  29.
The docking results of AutoDock Vina 12 guided by different pocket-finding algorithms for the Astex Diverse Set 29, consisting of 85 protein structures, are shown in Fig. 20. The corresponding RMSD distribution of the Top-1 predicted Vina poses is presented in Fig. 21, and the associated PLI scores are shown in Fig. 22.
Several striking features emerge from these graphs. First, although each individual Run of our model outperforms the docking accuracy achieved when guiding AutoDock Vina 12 using PUResNet 37 or Kalasanty 38, ensembling these individual Runs does not improve the overall docking accuracy. This happens even though the ensembled version of RAPID-Net achieves a higher ligand-pocket intersection, as shown in Fig. 22 and Table II.
Furthermore, unlike the PoseBusters 28 dataset, when AutoDock Vina 12 is guided by Kalasanty 38 on the Astex Diverse Set 29, it achieves better docking accuracy than when guided by PUResNet 37, even though PUResNet 37 has more structures with at least one true ligand binding pose completely within the search grid (72 vs. 70, respectively).
@@
Table II: Comparison of pocket predictions across different methods for Astex Diverse Set 29 consisting of 85 protein structures.
Figure 25: Ranks of subleading Vina poses that pass the RMSD $<$ 2 Å test when the Top-1 Vina pose fails for Astex Diverse Set 29.
Figure 26: In the 1G9V protein structure from Astex Diverse Set 29, two true ligand binding poses are observed within the same majority-voted pocket predicted by RAPID-Net. While the Top-1 Vina pose fails the RMSD test, the subleading one passes all validation tests.
Nevertheless, when considering ensemble accuracy–achieved either by the Top-1 pose or any subleading pose, as illustrated in Fig. 23–the ensembled version of RAPID-Net outperforms each of its individual Runs: 95.3% versus 92.9%, 87.1%, 89.4%, 90.6%, and 91.8%, respectively. Additionally, PUResNet 37 and Kalasanty 38 both achieve the same ensemble accuracy of 78.8%. Fig. 25 shows the distribution of subleading poses’ ranks passing the RMSD tests when the Top-1 Vina pose fails. For example, Fig. 26 illustrates the 1G9V protein structure, where a subleading pose ranked 39 – very far from the Top-1 pose – successfully passes all validation tests, while the Top-1 Vina pose fails.
A comparison of Tables I and II alongside Figs. 12 and 23 shows that RAPID-Net achieves a higher PLI rate on the Astex Diverse Set 29 than on the PoseBusters 28 dataset. However, the Top-1 docking accuracy is higher for PoseBusters 28, while the “at least one correct pose in the ensemble” accuracy is better for the Astex Diverse Set 29. These results further emphasize and illustrate that, in addition to accurate pocket identification, another major bottleneck in the docking process is the accurate reranking of the generated poses.
@@
VIII Evaluation on Coach420 and BU48 Datasets
For direct comparison with Kalasanty 38 and PUResNet 37, we evaluate the corresponding PLI rates on the Coach420 30 and BU48 31 datasets. Following 37, we exclude protein structures present in the training sets, resulting in 298 and 62 protein-ligand structures for Coach420 30 and BU48 31, respectively.
@@ VIII Evaluation on Coach420 and BU48 Datasets
Table III: Comparison of pocket predictions across different methods for Coach420 30 dataset consisting of 298 protein structures.
@@
Table IV: Comparison of pocket predictions across different methods for BU48 31 dataset consisting of 62 protein structures.
Figure 27: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 37, and Kalasanty 38 for the Coach420 30 dataset.
Figure 28: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 37, and Kalasanty 38 for the BU48 31 dataset.
Unlike the original PUResNet 37 paper where PLI values were evaluated only on protein structures where the distance-center-to-center (DCC) between predicted pocket centers and ligand centers was $\leq 4\text{\AA}$, we report the results for all protein structures, over the whole dataset. This comparison seems more appropriate because a rough initial approximation to set up the search grid is often sufficient to achieve successful docking, as illustrated in Fig. 6. Furthermore, as discussed in the next Section, the RAPID-Net model sometimes predicts meaningful “tunnels” or “bridges” to distant binding sites that indirectly influence ligand binding. In such cases, the DCC metric becomes irrelevant since it does not reflect the functional importance of these remote interactions.
Table III and Fig. 27 summarize the results for the Coach420 dataset 30. Except for Run 3, every RAPID-Net run surpasses PUResNet 37 and Kalasanty 38 in terms of an average PLI score. Furthermore, when comparing the number of protein structures in which at least one true ligand binding pose is entirely within the search grid, up to the largest grid with a threshold of $15\ \text{\AA}$, all RAPID-Net Runs outperform both PUResNet 37 and Kalasanty 38. Similarly, Table IV and Fig. 28 present the results for the BU48 dataset 31. Here, all RAPID-Net Runs, except Run 5, outperform both Kalasanty and PUResNet in terms of an average PLI and the number of protein structures where at least one true ligand binding pose is completely contained within the search grids.
Different RAPID-Net Runs have varying performances across the four test datasets, reflecting the inherent element of randomness in the model training. For example, on the PoseBusters 28 dataset, Run 2 has the fewest protein structures with viable predicted search grids, as shown in Table I. In contrast, Run 3 produces the fewest viable grids for the Astex Diverse Set 29 and Coach420 30, as shown in Tables II and III, while Run 5 shows the lowest PLI for BU48 31, as shown in Table IV. Notably, despite having lower coverage on Coach420 30 and BU48 31, Run 5 achieves the highest “at least one correct pose in the ensemble” rate on PoseBusters 28, as shown in Figure 12.
Figure 29: Thrombin (RCSB PDB: 1DWC). Pockets predicted by Kalasanty 38 and PUResNet 37 are shown as green and orange dots respectively. The majority-voted pocket predicted by our model is shown as cyan sticks where one can observe a distinct bulge towards the residues belonging to Exosite I.
@@
(b) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). PUResNet 37 does not predict any pockets for this protein structure, but Kalasanty 38 predicts one around the main binding site. One majority-voted pocket predicted by our model covers the active site and the other one is located near the histidine cluster responsible for the proton shuttle.
@@
(d) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). The interacting residues predicted by PUResNet V2 34 are located only near the active site, with secondary binding sites not being predicted.
Figure 30: Pockets predicted by Kalasanty 38 and PUResNet 37 are shown as green and orange dots, respectively. Majority-voted pockets predicted by RAPID-Net are represented by cyan sticks, while the minimally-reported pockets are shown by purple sticks. The interacting residues predicted by PUResNet V2 34 are shown in yellow.
Figure 31: Ls-AChBP (RCSB PDB: 2ZJV). The majority-voted pocket predicted by our model has a distinct bridge towards the Thr155-Asp160 region belonging to the F loop. Kalasanty 38 and PUResNet V1 37 do not predict any pockets for this protein structure, while PUResNet V2 34 predicts interacting residues only in the immediate vicinity of the binding pose.
These results further emphasize the importance of ensembling model predictions to capture all potential binding sites and account for variability in model training. By combining five RAPID-Net models, we mitigate performance defects of individual model Runs, yielding more robust and reliable results. At the same time, as can be observed from all the data that we presented, across all four test datasets, RAPID-Net consistently exhibits stronger generalization ability than both PUResNet 37 and Kalasanty 38.
@@
IX Identification of Allosteric Sites, Exosites, and Flexible Regions for Drug Design
To illustrate RAPID-Net’s ability to identify remote sites of therapeutic interest, we consider four proteins where such sites are well documented.
Thrombin (RCSB PDB: 1DWC) The RCSB PDB entry 1DWC 65 is the crystal structure of human $\alpha$-thrombin in complex with the inhibitor MD-805. Thrombin plays a pivotal role in the coagulation cascade and is therefore a key target in treating acute coronary syndromes 66. As shown in Fig. 29, Kalasanty 38 and PUResNet 37 predict binding pockets largely surrounding the ligand. By contrast, our model, RAPID-Net, detects an additional bulge extending towards residues 71, 73, 75, 76, and 77. These residues belong to the anion-binding Exosite I, which interacts with negatively charged substrates and cofactors such as fibrinogen, thrombomodulin, and COOH-terminal peptide of hirudin 67; 68; 69; 70; 71. Our model’s prediction, a “tunnel” connecting the active site to the Exosite I in Fig. 29, suggests possible long-range interactions, consistent with previous studies of long-range allosteric communication in thrombin 71.
Human Carbonic Anhydrase I (RCSB PDB: 1AZM) and II (RCSB PDB: 3HS4) Human carbonic anhydrase I (hCAI) in complex with a sulfonamide drug (PDB: 1AZM) 72 and hCAII in complex with acetazolamide (AZM) (PDB: 3HS4) 73 are key targets for glaucoma treatment 74 and diuretic therapy 75, with broader potential for treating obesity, cancer, and Alzheimer’s disease 75. In both isoforms, His64 acts as the primary proton shuttle 76; 77; 78.
As shown in Fig. 30(a), while PUResNet V1 37 and Kalasanty 38 predict the primary ligand binding region in 1AZM, our model additionally predicts a hook-shaped bulge wrapping around His64, suggesting an allosteric interaction.
In hCAII (3HS4), a histidine cluster (His3, His4, His10, His15, His17, His64) mediates proton exchange between the active site and the environment 75. Notably, 3HS4 contains two additional AZM molecules in shallow surface pockets–one previously identified 79; 80 and another in a novel binding site 73.
@@ IX Identification of Allosteric Sites, Exosites, and Flexible Regions for Drug Design
In comparison, for this protein structure, Kalasanty 38 predicts only the main binding site, PUResNet V1 37 detects none, and PUResNet V2 34 identifies only the interacting residues within the main binding site.
Ls-AChBP (RCSB PDB: 2ZJV). RCSB PDB ID 2ZJV 81 represents Lymnaea stagnalis acetylcholine-binding protein (AChBP) bound to neonicotinoid clothianidin. AChBP serves as a model for nicotinic acetylcholine receptors (nAChRs) and their allosteric transitions 82, providing insights into therapies for neurological disorders such as Alzheimer’s disease, schizophrenia, depression, attention deficit hyperactivity disorder, and tobacco addiction 83; 84; 85.
For the 2ZJV protein structure, neither Kalasanty 38 nor PUResNet V1 37 predict any pockets. However, the majority-voted pockets from our model and the interacting residues of PUResNet V2 34 illustrated in Fig. 31 are predicted around all five true ligand binding positions.
Furthermore, our model’s pockets include distinct bridges toward residues Thr155-Asp160 in the flexible F-loop, suggesting indirect ligand interactions. This finding aligns with the results from 81, who highlighted that induced-fit movements of loop regions, including the F-loop, are essential for ligand recognition by neonicotinoids.
@@ X Discussion and Conclusions
When guided by RAPID-Net, AutoDock Vina 12 outperforms DiffBindFR 43 by over 5% in blind docking accuracy on the PoseBusters 28 dataset, highlighting the direct relationship between improved pocket identification and increased docking accuracy. Furthermore, RAPID-Net provides precise and compact search grids, enabling the docking of ligands to large proteins such as 8F4J from the PoseBusters 28 dataset, which tools like AlphaFold 3 33 cannot process in whole, highlighting the critical importance of accurate and focused search areas for effective docking.
We found that another major factor limiting docking accuracy is the proper reranking of generated poses. Notably, when assessing accuracy based on whether at least one pose in the ensemble is correct (instead of only the top-ranked pose), AutoDock Vina 12 surpasses AlphaFold 3 33 in accuracy on PoseBusters 28 dataset. This advantage becomes even more significant in blind docking scenarios. With RAPID-Net’s guidance, AutoDock Vina 1.2.5 12 exceeds AlphaFold 3’s performance by over 10% on the PoseBusters 28 dataset. These findings suggest that, in addition to precise pocket identification, developing improved reranking tools could further enhance the docking accuracy of our combined scheme by effectively selecting the most favorable poses from those generated.
@@
Finally, our model demonstrated the ability to identify “bridges” to distal sites located more than $6.5\ \text{\AA}$ away from the ligand binding site, which are crucial for therapeutic applications, making RAPID-Net a promising tool for uncovering such areas and providing valuable insights for broader therapeutic strategies.
In summary, RAPID-Net not only improves the accuracy of pocket identification and subsequent docking but also opens new avenues for targeting remote therapeutic sites, highlighting its potential in the field of computational drug discovery.

### v2

@@ VIII Evaluation on the Astex Diverse Set
Figure 20: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the Astex Diverse Set 33.
Figure 21: Docking accuracy of AutoDock Vina 12 for the Astex Diverse Set 33 when guided by different pocket prediction algorithms, comparing Top-1 and ensemble accuracy.
The docking results of AutoDock Vina 12 guided by different pocket-finding algorithms for the Astex Diverse Set 33, consisting of 85 protein structures, are shown in Fig. 18. The corresponding RMSD distribution of the Top-1 predicted Vina poses is presented in Fig. 19, and the associated PLI scores are shown in Fig. 20.
Several striking features emerge from these graphs. First, although each individual Run of our model outperforms the docking accuracy achieved when guiding AutoDock Vina 12 using PUResNet 69 or Kalasanty 70, ensembling these individual Runs does not improve the overall docking accuracy. This happens even though the ensembled version of RAPID-Net achieves a higher ligand-pocket intersection, as shown in Fig. 20 and Table II.
Furthermore, unlike the PoseBusters 28 dataset, when AutoDock Vina 12 is guided by Kalasanty 70 on the Astex Diverse Set 33, it achieves better docking accuracy than when guided by PUResNet 69, even though PUResNet 69 has more structures with at least one true ligand binding pose completely within the search grid (72 vs. 70, respectively).
@@
Table II: Comparison of pocket predictions across different methods for Astex Diverse Set 33 consisting of 85 protein structures.
Figure 22: In the 1G9V protein structure from Astex Diverse Set 33, two true ligand binding poses are observed within the same majority-voted pocket predicted by RAPID-Net. While the Top-1 Vina pose fails the RMSD test, the subleading one passes all validation tests.
Nevertheless, when considering sampling accuracy–achieved either by the Top-1 pose or any subleading pose, as illustrated in Fig. 21–the ensembled version of RAPID-Net outperforms each of its individual Runs: 95.3% versus 92.9%, 87.1%, 89.4%, 90.6%, and 91.8%, respectively. Additionally, PUResNet 69 and Kalasanty 70 both achieve the same sampling accuracy of 78.8%. A comparison of Tables I and II alongside Figs. 12 and 21 shows that RAPID-Net achieves a higher PLI rate on the Astex Diverse Set 33 than on the PoseBusters 28 dataset. However, the Top-1 docking accuracy is higher for PoseBusters 28, while the “at least one correct pose in the ensemble” accuracy is better for the Astex Diverse Set 33. These results further emphasize and illustrate that, in addition to accurate pocket identification, another major bottleneck in the docking process is the accurate reranking of the generated poses.
@@
IX Evaluation on Coach420 and BU48 Datasets
For direct comparison with Kalasanty 70 and PUResNet 69, we evaluate the corresponding PLI rates on the Coach420 34 and BU48 35 datasets. Following 69, we exclude protein structures present in the training sets, resulting in 298 and 62 protein-ligand structures for Coach420 34 and BU48 35, respectively.
@@ IX Evaluation on Coach420 and BU48 Datasets
Table III: Comparison of pocket predictions across different methods for Coach420 34 dataset consisting of 298 protein structures.
@@
Table IV: Comparison of pocket predictions across different methods for BU48 35 dataset consisting of 62 protein structures.
Figure 23: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the Coach420 34 dataset.
Figure 24: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the BU48 35 dataset.
Figure 25: For the Nsp12 protein 64, the pockets predicted by Kalasanty 70 and PUResNet 69 – shown as green and orange dots, respectively – and the likely interacting residues predicted by PUResNet V2 66 correspond to the orthosteric remdesivir binding site 64. In contrast, RAPID-Net not only identifies the orthosteric site – corresponding to the majority-voted pocket, shown as cyan sticks – but also successfully detects putative secondary and allosteric pockets reported in previous studies 98; 99, which correspond to the minimally-reported pockets predicted by our model, shown as purple sticks.
Figure 26: Thrombin (RCSB PDB: 1DWC). Unlike the pockets predicted by Kalasanty 70 and PUResNet 69 –shown as green and orange dots, respectively – the majority-voted pocket predicted by our model, shown as cyan sticks, has a distinct bulge extending toward residues associated with Exosite I.
@@
(b) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). PUResNet 69 does not predict any pockets for this protein structure, but Kalasanty 70 predicts one around the main binding site. One majority-voted pocket predicted by our model covers the active site and the other one is located near the histidine cluster responsible for the proton shuttle.
@@
(d) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). The interacting residues predicted by PUResNet V2 66 are located only near the active site, with secondary binding sites not being predicted.
Figure 27: Majority-voted pockets predicted by RAPID-Net are represented by cyan sticks, while the minimally-reported pockets are shown by purple sticks. The interacting residues predicted by PUResNet V2 66 are shown in yellow.
Figure 28: Ls-AChBP (RCSB PDB: 2ZJV). The majority-voted pocket predicted by our model has a distinct bridge towards the Thr155-Asp160 region belonging to the F loop. Kalasanty 70 and PUResNet V1 69 do not predict any pockets for this protein structure, while PUResNet V2 66 predicts interacting residues only in the immediate vicinity of the binding pose.
Unlike the original PUResNet 69 paper where PLI values were evaluated only on protein structures where the distance-center-to-center (DCC) between predicted pocket centers and ligand centers was $\leq 4\text{\AA}$, we report the results for all protein structures, over the whole dataset. This comparison seems more appropriate because a rough initial approximation to set up the search grid is often sufficient to achieve successful docking, as illustrated in Fig. 5. Furthermore, as discussed in the next Section, the RAPID-Net model sometimes predicts meaningful “tunnels” or “bridges” to distant binding sites that indirectly influence ligand binding. In such cases, the DCC metric becomes irrelevant since it does not reflect the functional importance of these remote interactions.
Table III and Fig. 23 summarize the results for the Coach420 dataset 34. Except for Run 3, every RAPID-Net run surpasses PUResNet 69 and Kalasanty 70 in terms of an average PLI score. Furthermore, when comparing the number of protein structures in which at least one true ligand binding pose is entirely within the search grid, up to the largest grid with a threshold of $15\ \text{\AA}$, all RAPID-Net Runs outperform both PUResNet 69 and Kalasanty 70. Similarly, Table IV and Fig. 24 present the results for the BU48 dataset 35. Here, all RAPID-Net Runs, except Run 5, outperform both Kalasanty and PUResNet in terms of an average PLI and the number of protein structures where at least one true ligand binding pose is completely contained within the search grids.
Different RAPID-Net Runs have varying performances across the four test datasets, reflecting the inherent element of randomness in the model training. For example, on the PoseBusters 28 dataset, Run 2 has the fewest protein structures with viable predicted search grids, as shown in Table I. In contrast, Run 3 produces the fewest viable grids for the Astex Diverse Set 33 and Coach420 34, as shown in Tables II and III, while Run 5 shows the lowest PLI for BU48 35, as shown in Table IV. Notably, despite having lower coverage on Coach420 34 and BU48 35, Run 5 achieves the highest “at least one correct pose in the ensemble” rate on PoseBusters 28, as shown in Fig. 12.
These results further emphasize the importance of ensembling model predictions to capture all potential binding sites and account for variability in model training. By combining five RAPID-Net models, we mitigate performance defects of individual model Runs, yielding more robust and reliable results. At the same time, as can be observed from all the data that we presented, across all four test datasets, RAPID-Net consistently exhibits stronger generalization ability than both PUResNet 69 and Kalasanty 70.
@@
X Identification of Allosteric Sites, Exosites, and Flexible Regions for Drug Design
As noted earlier, a common problem with Deep Learning-based pocket prediction methods is that they may miss secondary pockets due to the scarcity of labeled training data, while classical methods may produce “blurry” pockets 65. To demonstrate the ability of RAPID-Net to accurately detect both primary and secondary pockets, we focus on the therapeutically relevant case of the Nsp12 protein, an RNA-dependent RNA polymerase (RdRp), which plays a central role in viral RNA replication and is a key target for the development of antiviral drugs against SARS-CoV-2.
RNA-dependent RNA Polymerase (RCSB PDB: 7BV2). As illustrated in Fig. 25, RAPID-Net successfully identifies the orthosteric remdesivir binding site 64 – containing residues 448, 545, 553, 555, 618, 759, 760, and 761 – as its “majority-voted” pocket, consistent with predictions from other pocket prediction methods, PUResNet V1 69, Kalasanty 70, and PUResNet V2 66. In addition to the main binding site, computational studies have proposed allosteric 98 and secondary pockets 99 that can be used for drug development. Unlike other pocket predictors, RAPID-Net also highlights putative allosteric and secondary pockets reported in the literature 98; 99, demonstrating its ability to recover both primary and secondary binding sites.
Additionally, to illustrate RAPID-Net’s ability to identify remote sites of therapeutic interest, we consider four proteins where such sites are well documented.
Thrombin (RCSB PDB: 1DWC). The RCSB PDB entry 1DWC 100 is the crystal structure of human $\alpha$-thrombin in complex with the inhibitor MD-805. Thrombin plays a pivotal role in the coagulation cascade and is therefore a key target in treating acute coronary syndromes 101. As shown in Fig. 26, Kalasanty 70 and PUResNet 69 predict binding pockets largely surrounding the ligand. By contrast, our model, RAPID-Net, detects an additional bulge extending towards residues 71, 73, 75, 76, and 77. These residues belong to the anion-binding Exosite I, which interacts with negatively charged substrates and cofactors such as fibrinogen, thrombomodulin, and COOH-terminal peptide of hirudin 102; 103; 104; 105; 106. Our model’s prediction, a “tunnel” connecting the active site to the Exosite I in Fig. 26, suggests possible long-range interactions, consistent with previous studies of long-range allosteric communication in thrombin 106.
Human Carbonic Anhydrase I (RCSB PDB: 1AZM) and II (RCSB PDB: 3HS4). Human carbonic anhydrase I (hCAI) in complex with a sulfonamide drug (PDB: 1AZM) 107 and hCAII in complex with acetazolamide (AZM) (PDB: 3HS4) 108 are key targets for glaucoma treatment 109 and diuretic therapy 110, with broader potential for treating obesity, cancer, and Alzheimer’s disease 110. In both isoforms, His64 acts as the primary proton shuttle 111; 112; 113.
As shown in Fig. 27(a), while PUResNet V1 69 and Kalasanty 70 predict the primary ligand binding region in 1AZM, our model additionally predicts a hook-shaped bulge wrapping around His64, suggesting an allosteric interaction.
In hCAII (3HS4), a histidine cluster (His3, His4, His10, His15, His17, His64) mediates proton exchange between the active site and the environment 110. Notably, 3HS4 contains two additional AZM molecules in shallow surface pockets–one previously identified 114; 115 and another in a novel binding site 108.
@@ X Identification of Allosteric Sites, Exosites, and Flexible Regions for Drug Design
In comparison, for this protein structure, Kalasanty 70 predicts only the main binding site, PUResNet V1 69 detects none, and PUResNet V2 66 identifies only the interacting residues within the main binding site.
Ls-AChBP (RCSB PDB: 2ZJV). RCSB PDB ID 2ZJV 116 represents Lymnaea stagnalis acetylcholine-binding protein (AChBP) bound to neonicotinoid clothianidin. AChBP serves as a model for nicotinic acetylcholine receptors (nAChRs) and their allosteric transitions 117, providing insights into therapies for neurological disorders such as Alzheimer’s disease, schizophrenia, depression, attention deficit hyperactivity disorder, and tobacco addiction 118; 119; 120.
For the 2ZJV protein structure, neither Kalasanty 70 nor PUResNet V1 69 predict any pockets. However, the majority-voted pockets from our model and the interacting residues of PUResNet V2 66 illustrated in Fig. 28 are predicted around all five true ligand binding positions.
Furthermore, our model’s pockets include distinct bridges toward residues Thr155-Asp160 in the flexible F-loop, suggesting indirect ligand interactions. This finding aligns with the results from 116, who highlighted that induced-fit movements of loop regions, including the F-loop, are essential for ligand recognition by neonicotinoids.
@@ XI Discussion and Conclusions
When guided by RAPID-Net, AutoDock Vina 12 outperforms DiffBindFR 27 by over 5% in blind docking accuracy on the PoseBusters 28 dataset, highlighting the direct relationship between improved pocket identification and increased docking accuracy. Furthermore, RAPID-Net provides precise and compact search grids, enabling the docking of ligands to large proteins such as 8F4J from the PoseBusters 28 dataset, which tools like AlphaFold 3 30 cannot process in whole, highlighting the critical importance of accurate and focused search areas for cost-effective docking.
We found that another major factor limiting docking accuracy is the reranking of generated poses. In addition to precise pocket identification, developing improved reranking tools could further enhance the overall performance of our combined scheme by more effectively selecting the most favorable poses from the generated ensemble.
@@
Furthermore, RAPID-Net demonstrates the ability to identify “bridges” to distal sites further than $6.5\ \text{\AA}$ from the primary pocket that are often critical for therapeutic intervention. This makes RAPID-Net a promising tool for targeting remote therapeutic sites and broadening the scope of structure-based drug design.
Additionally, beyond improved docking accuracy, in our future work, we plan to integrate RAPID-Net into de novo pocket-conditioned drug design pipelines by guiding the generation of novel compounds based on the shape features of the predicted pockets. Although our current study does not directly exploit these features, using the predicted region solely as a search box, as illustrated in Fig. 4, they may prove highly valuable for generative models 121; 122; 123; 124; 125.